Minute of KAGRA commissioning meeting
- Date: 2018/4/17 11AM~12PM
- Attendance: Miyakawa, Kokeyama, Aso, Uchiyama, Miyoki, Akutsu
- eZuce: Yamamoto, Suzuki, Shoda, Nagano
1. Report of work of the last week and plan of work
- Door was closed and pump down started. One large leak was found and it is being searched where it is.
- IFO, ETMs: Enomono, Ushiba
- Michelson fringe was seen on DC and RF, but still mirror motion is too large.
- Needs more work for damping for both ETMs, L2P of BS.
- IO chassis of k1ioo was disconnected somehow. It is being checked now.
- Recoding minute_raw of k1fw1 moved to SSD at k1tw1. All fw0,1 tw0,1 are working well without any gap on data.
- Noise injection needs some preparations for speakers.
- Schnupp asymmetry need some preparations.
- Other plan seems to be ready.
- CBC will perform an online analysis.
3. Tasks or discussions
Selection of full channel -> DGS Yamamoto
Selection of science channel -> Commissioning Izumi
- Screens at control room
- Needs laser safety information.
GW alert: online alert in the control room might be prepared. -> Togoshi
- Laser power of oplev started dropping, specially lasers where are close to cryostat for year term.
- Phase 1 summarize should talk wide range view.
- Type A has double stable point.
- Monitor for water level, and vacuum gage for weekend should be implemented.
- Patrol light should be implemented soon, even manual one.
- Next meeting: 4/24